Grade 11Mechanics

Balls are thrown vertically upwards in such a way that next ball is thrown when previous one is at maximum height if maximum height is5m find number of balls thrown in a minute

Bhawika
10 Years agoGrade 11

1 Answer

Paras Verma
10 Years ago
Initial velocity-
v2=u2+2as
02=u2-2(10)5
u=10
v=u+at
0=10-10t
t=1s
Therefore, in a minute you’ll be able to throw 60 balls.